# On the Role of Entity and Event Level Conceptualization in Generalizable Reasoning: A Survey of Tasks, Methods, Applications, and Future Directions

### v1

@@ Abstract
Entity- and event-level conceptualization, as fundamental elements of human cognition, plays a pivotal role in generalizable reasoning. This process involves abstracting specific instances into higher-level concepts and forming abstract knowledge that can be applied in unfamiliar or novel situations, which can enhance models’ inferential capabilities and support the effective transfer of knowledge across various domains. Despite its significance, there is currently a lack of a systematic overview that comprehensively examines existing works in the definition, execution, and application of conceptualization to enhance reasoning tasks. In this paper, we address this gap by presenting the first comprehensive survey of 150+ papers, categorizing various definitions, resources, methods, and downstream applications related to conceptualization into a unified taxonomy, with a focus on the entity and event levels. Furthermore, we shed light on potential future directions in this field and hope to garner more attention from the community.
@@ 1 Introduction
Concepts are the glue that holds our mental world together.”– Murphy (2004)
@@
In general terms, conceptualization refers to the process of consolidating specific instances with shared properties or characteristics into a cohesive concept that represents a vast collection of instances. It is a sub-type of abstraction Giunchiglia and Walsh (1992), but specifically requires the presence of a concept as the base for such abstraction. With proper conceptualization, abstract knowledge can be subsequently derived by associating original knowledge at the instance level with that concept. When encountering unfamiliar or novel scenarios, concepts in abstract knowledge can be instantiated to new instances to support downstream reasoning Tenenbaum et al. (2011). This process can occur at various levels, including entity Wu et al. (2012); Liang et al. (2017); Alukaev et al. (2023); Liu et al. (2023c), event He et al. (2024); Wang et al. (2024a, c), paragraph/document Falke and Gurevych (2019); Falke et al. (2017), and system levels Subramonian et al. (2023); Kadioglu and Kleynhans (2024), ultimately forming a hierarchy that contribute to a comprehensive understanding and representation of knowledge.
@@
Despite its significance, the field lacks a comprehensive and unified taxonomy to categorize existing research on conceptualization. This has led to several drawbacks, such as the inconsistent use of the term “conceptualization” across different studies, resulting in varying definitions despite a common underlying meaning. Additionally, the methods for conceptualizing different types of instances in a scalable and accurate manner remain unclear. Finally, it is essential to summarize the benefits that conceptualization can bring to downstream tasks to gather insights for future applications and new research directions.
@@
To address these issues, we present the first-ever survey that systematically taxonomizes conceptualization. Firstly, in Section 2, we present a hierarchical definition of conceptualization based on different semantic levels of instances being conceptualized, namely: entity, event, paragraph/document, and system. In later sections, we focus on two main types of conceptualization based on the entity and event levels, as they are most prevalent in existing literature and play a key role in human reasoning. We review more than 150 papers and organize them into four main categories, as shown in Figure 2. We summarize the main representative tasks and datasets available for these types of conceptualization in Section 3. Subsequently, in Section 4, we categorize conceptualization acquisition methods into extraction, retrieval, and generative-based methods. The downstream benefits of conceptualization are discussed in Section 5, with a specific focus on several reasoning tasks. Finally, in Section 6, we propose two future directions that can be benefited from conceptualization. We hope our work can serve as a practical handbook for researchers and pave the way for further advancements in conceptualization.
2 The Hierarchy of Conceptualization
@@
We first define the hierarchy of conceptualization according to the type of instances being conceptualized. They are categorized into four levels: entity level, event level, document level, and system level. Running examples are shown in Figure 1.
@@
Reasoning (§5.2) [ Fang et al. (2024); Zheng et al. (2023); Cui et al. (2017); Ou et al. (2008) , leaf, text width=47em ] ] [ Others (§5.3) [ Choi et al. (2018); Li et al. (2022); Dai and Zeng (2023); Jiang et al. (2023); Li et al. (2023); Feng et al. (2023a)
@@
Entity-level conceptualization involves grouping multiple entities under a shared concept Yang et al. (2021); Peng et al. (2022). It is the most common form of conceptualization in human cognition and is frequently applied for knowledge acquisition Carey (1991); Murphy (2004). For instance, entities like “apple,” “pear,” and “grape,” can be categorized together under the broader concept of “fruit.” By doing so, abstract knowledge can be derived by reintegrating the concept into the context of specific instances, such as the assertion “fruit is delicious,” with “apple is delicious” serving as the specific source. When someone encounters an unknown fruit, they can quickly understand its properties by associating it with the abstract knowledge of fruit, such as its possible taste or nutrition.
Event Level:
@@
While a concept can capture the semantic meaning of a group of entities, it can also represent events at a higher level of conceptualization. Event-level conceptualization aims to broaden the scope from entities to include events as well He et al. (2024); Wang et al. (2024c). It seeks to associate different events under a shared concept that preserves the original semantic meaning to the maximum extent possible. For instance, activities like “Sam playing with his dog,” “Alex dancing in the club,” and “Bob doing yoga” can all be conceptualized as “relaxing events.” Abstract knowledge can then follow, stating that “If someone engages in relaxing events, they feel happy and relaxed.” When someone encounters an unknown or unfamiliar event, such as “Charlie likes painting the sunset,” they can infer that painting the sunset is a relaxing event and that Charlie feels happy and relaxed when doing so.
Document Level:
Document-level conceptualization further extends the scope of the instance from entities and events to paragraphs or even entire documents. One representative task in this category is abstractive summarization Ladhak et al. (2022); Wang et al. (2019); Lin and Ng (2019), where the objective is to generate a summary that captures the main ideas and essential information while maintaining the overall meaning and context of the original text. For example, documents on topics like “A Study on Climate Change,” “The Impact of Global Warming,” and “The Future of Renewable Energy” can all be abstracted under the concept of “Environmental Studies.” Abstract knowledge can then be derived, such as “Environmental studies discuss the impact of human activities on the environment and potential solutions.” This process is challenging because it requires a delicate balance between preserving essential details related to the concept and reducing the length of the text.
System Level:
@@
Finally, system-level conceptualization aims to simplify the understanding of a complex system by abstracting its behavior and functionality into a higher-level representation. There is no fixed definition of system-level conceptualization, as it can vary depending on the context and the system being considered. A representative example in NLP is recent work by Subramonian et al. (2023), where the authors provide a systematic categorization of NLP tasks based on their objectives and characteristics while neglecting the detailed format of input/output and the dataset the tasks are evaluated on. Abstract knowledge typically comes in the form of knowledge or facts associated with the derived conceptualization, which may vary depending on the context.
@@ 3 Tasks and Datasets
While all levels of conceptualization play a pivotal role in knowledge representation, those at the entity and event levels are the most fundamental due to their unique importance in human cognition and generalizable reasoning. Therefore, in later sections, we specifically focus on entity and event level conceptualizations. We first discuss existing literature on the concept linking task and examine currently available resources for these two types of conceptualizations. Statistical comparisons between different resources are shown in Table 1. For datasets that also serve as evaluation benchmarks, we mark their associated tasks with classification task (CLS) and generation task (GEN).
3.1 Concept Linking Tasks
The main task of conceptualization can be formulated as a concept linking task, where the goal is to link an instance $i$ to a concept $c$ such that $i$ can be semantically represented by $c$. It is challenging due to the infinite number of possible instance-concept pairs. Previous approaches, such as those byBrauer et al. (2010); Yates et al. (2015), have attempted to further restrict the task to linking instances to a limited set of strict ontologies using heuristic or statistical methods. The task can also be formulated with a generative objective, which requires a model to generate $c$ directly given $i$ as input.
3.2 Entity-level Conceptualization
@@
3.3 Event-level Conceptualization
@@ 4 Conceptualization Acquisition Methods
In this section, we discuss methods for performing or collecting entity and event-level conceptualizations. We categorize them into three types: extraction, retrieval, and generative-based methods, which are briefly demonstrated in Figure 3. We provide more discussions in Appendix A.
@@ 4.2.2 Neural-Based Retrieval
Neural-based retrieval methods overcome previous limitations by leveraging neural networks (or language models) to learn the semantic representations of the input instance and the concepts in the knowledge base or concept taxonomy. Then, the similarity between the input instance and the concepts can be calculated based on the learned representation embeddings. This approach can be benefitted by the advancement in language modeling, such as BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), and DeBERTa He et al. (2021, 2023). The most representative work in neural-based concept retrieval is AbstractATOMIC He et al. (2024). It uses GlossBERT Huang et al. (2019) to encode concepts (from WordNet and Probase) and instances (extracted from events in ATOMIC Sap et al. (2019)) into embeddings and leverage cosine similarity and human annotations to collect conceptualizations in a large scale manner. Other methods Wang et al. (2024c); Zhang et al. (2020, 2022); Lu et al. (2023); Wang et al. (2023b); Gao et al. (2022); Becker et al. (2021) similarly adopt different strategies in leveraging LMs as encoders, expanding the coverage of instances,training retrieval models. Despite their promising results, these methods are limited by their need for extensive labeled data, reliance on the completeness and accuracy of the knowledge base, and inability to retrieval new concepts that are out of training data.
@@ 4.3.1 Fine-Tuning-Based Generative Methods
Fine-tuning-based generative methods aim to take an entity or event as input and generate the concept directly via a fine-tuned generative language model. This approach allows the model to generate conceptualizations for new instances and offers maximum flexibility of the input. Several methods Peng et al. (2022); Yuan et al. (2023); He et al. (2024); Wang et al. (2024c, b, 2023b) have adopted this paradigm in training generative conceptualizers, based on models such as GPT2 Radford et al. (2019), BART Lewis et al. (2020), and T5 Raffel et al. (2020), for automated conceptualization acquisition. These methods typically train LMs on human-annotated or pre-existing conceptualization resources and yield outstanding results. However, fine-tuning-based generative methods are limited by their high computational cost, time-consuming and resource-intensive data collection, uncertain performance across diverse domains, and relatively low quality of novel concepts compared to human annotations. While these are common limitations associated with fine-tuned generative models, zero-shot generative methods using powerful LLMs and advanced prompting techniques potentially address these issues.
@@ 5 Downstream Applications
In this section, we discuss downstream tasks that can be benefited from applying conceptualizations. An overview of performances by different methods that leverage conceptualization, evaluated on various benchmarks, are shown in Figure 4.
Figure 4: Empirical benefits of conceptualization in methods across various benchmarks. All reported results are sourced from respective original papers.
@@ 5.1 Commonsense Reasoning
The task of generative commonsense inference modeling (COMET; Bosselut et al. (2019); Hwang et al. (2021)) aims to complete an inferential commonsense knowledge given a head event and a commonsense relation. State-of-the-art methods for COMET mainly fine-tune language models on large-scale commonsense knowledge bases, which suffer from data sparsity and lack of diversity in commonsense knowledge. Although transfer from LLMs helps West et al. (2022, 2023), distilled knowledge tends to be too easy for models to learn and converge to trivial inferences. To address these issues, Wang et al. (2023b) proposed to leverage conceptualization as knowledge augmentation tools to improve COMET. Conceptualizations are first derived from head events to obtain abstracted events. Then, the tail of the original commonsense knowledge is placed back to the abstracted event to form abstracted commonsense knowledge. These derived abstract knowledge are then integrated with the original knowledge in commonsense knowledge bases to enrich the diversity of commonsense knowledge. Experiments results show consistent improvement in models’ performances. Wang et al. (2024a) further show that, by instantiating conceptualizations in abstract knowledge back to other novel instances, models can be further improved by training with newly instantiated knowledge. Liu et al. (2023a) also proposed a task that aims to generate diverse sentences describing concept relationships in various everyday scenarios. Conceptualizations and associated abstract knowledge can further boost models’ performances on this task.
@@ 6 Future Directions and Conclusions
6.1 Controllable Generation and Hallucination Reduction
@@ 6.2 Modeling Changes in Distribution
Conceptualization also plays a pivotal role in building reasoning systems that can capture situational changes in distribution to achieve System II reasoning Sloman (1996); Kahneman (2011). Among the several components that make up System II reasoning, a key element is the ability to reason with situational changes in distribution Bengio et al. (2021, 2019). These changes are triggered by environmental factors and actions by the agents themselves or others, especially when dealing with non-stationarities  Bengio (2017). This ability can be achieved by dynamically recombining existing concepts in the given environment or action and learning from the resultant situational changes Lake and Baroni (2018); Bahdanau et al. (2019); de Vries et al. (2019). For instance, consider the event “PersonX is driving a car on a sunny day.” A change in the weather from sunny to rainy could cause a different outcome, such as “PersonX becomes more cautious and drives slower.” This illustrates that a change in weather conditions can lead to a change in the driver’s behavior, representing an environmental change that triggers situational changes within the distribution of different weather conditions. In this process, the model is required to infer different changes that can possibly occur within a single event as the context, and reason about the potential outcome of each change. To model the distribution of different changes within an event, conceptualization can be used to represent the different states of the environment or action Wang and Song (2024). The model can then reason about the changes in distribution by manipulating the granularity of conceptualized changes. This type of distributional conceptualization not only provides an ontology for modeling the distribution of different changes within an event, but also assists the model in reasoning about the potential outcomes with appropriate abstract knowledge. Future works can leverage LLMs to curate benchmark datasets via sequential conceptualization generation and develop advanced systems for System II reasoning.

### v2

@@ Abstract
Conceptualization, a fundamental element of human cognition, plays a pivotal role in human generalizable reasoning. Generally speaking, it refers to the process of sequentially abstracting specific instances into higher-level concepts and then forming abstract knowledge that can be applied in unfamiliar or novel situations. This enhances models’ inferential capabilities and supports the effective transfer of knowledge across various domains. Despite its significance, the broad nature of this term has led to inconsistencies in understanding conceptualization across various works, as there exists different types of instances that can be abstracted in a wide variety of ways. There is also a lack of a systematic overview that comprehensively examines existing works on the definition, execution, and application of conceptualization to enhance reasoning tasks. In this paper, we address these gaps by first proposing a categorization of different types of conceptualizations into four levels based on the types of instances being conceptualized, in order to clarify the term and define the scope of our work. Then, we present the first comprehensive survey of over 150 papers, surveying various definitions, resources, methods, and downstream applications related to conceptualization into a unified taxonomy, with a focus on the entity and event levels. Furthermore, we shed light on potential future directions in this field and hope to garner more attention from the community.
@@ 1 Introduction
In general terms, conceptualization refers to the process of consolidating specific instances with shared properties or characteristics into a cohesive concept that represents a vast collection of instances. It is a sub-type of abstraction Giunchiglia and Walsh (1992), but specifically requires the presence of a concept as the base for such abstraction. With proper conceptualization, abstract knowledge can be subsequently derived by associating original knowledge at the instance level with that concept. When encountering unfamiliar or novel scenarios, concepts in abstract knowledge can be instantiated to new instances to support downstream reasoning Tenenbaum et al. (2011). This process can occur at various levels, including entity Wu et al. (2012); Liang et al. (2017); Alukaev et al. (2023); Liu et al. (2023c); Zhang et al. (2025), event He et al. (2024); Wang et al. (2024a, c), paragraph/document Falke and Gurevych (2019); Falke et al. (2017), and system levels Subramonian et al. (2023); Kadioglu and Kleynhans (2024), ultimately forming a hierarchy that contribute to a comprehensive understanding and representation of knowledge.
@@
Despite its significance, the field lacks a comprehensive and unified taxonomy to categorize existing research on conceptualization. On the one hand, the term “conceptualization” is inherently broad, encompassing various types of conceptualizations across different instances and performed in various ways, all included under a single term. As illustrated in Figure 1, the conceptualization of entities and documents requires two distinct paradigms; however, the current terminology fails to adequately address these differences. This has led to confusion and miscommunication among works that apply conceptualization in their methodologies. On the other hand, the methods for conceptualizing different types of instances in a scalable and accurate manner remain unclear. Finally, it is essential to summarize the benefits that conceptualization can bring to downstream tasks to gather insights for future applications and new research directions.
@@
To address these issues, we present the first-ever survey that systematically taxonomizes conceptualization. First, in Section 2, we define four types of conceptualization based on different semantic levels of the instances being conceptualized: entity, event, document, and system. In later sections, we focus on two main types of conceptualization based on the entity and event levels, as they are largely uncharted in existing literature and play a key role in human reasoning. We then propose a set of objectives to select and survey papers that feature conceptualization as their core idea, review more than 150 papers, and organize them into three main categories, as shown in Figure 2. We summarize the main representative tasks and datasets available for these types of conceptualization in Section 3. Subsequently, in Section 4, we categorize conceptualization acquisition methods into extraction, retrieval, and generative-based methods. The downstream benefits of conceptualization are discussed in Section 5, with a specific focus on several reasoning tasks. Finally, in Section 6, we propose two future directions that can benefit from conceptualization. We hope our work can serve as a practical handbook for researchers and pave the way for further advancements in the field of conceptualization.
2 Four Levels of Conceptualization
@@
We first define four levels of conceptualization according to the type of instances being conceptualized. They are categorized into four levels: entity level, event level, document level, and system level. Running examples are shown in Figure 1.
@@
Reasoning (§5.2) [ Fang et al. (2024); Zheng et al. (2023); Cui et al. (2017); Ou et al. (2008); Ding et al. (2024); Xu et al. (2024) , leaf, text width=47em ] ] [ Others (§5.3) [ Choi et al. (2018); Li et al. (2022); Dai and Zeng (2023); Jiang et al. (2023); Li et al. (2023); Feng et al. (2023a)
@@
Entity Level: Entity-level conceptualization involves grouping multiple entities under a shared concept Yang et al. (2021); Peng et al. (2022). It is the most common form of conceptualization in human cognition and is frequently applied for knowledge acquisition Carey (1991); Murphy (2004). For instance, entities like “apple,” “pear,” and “grape,” can be categorized together under the broader concept of “fruit.” By doing so, abstract knowledge can be derived by reintegrating the concept into the context of specific instances, such as the assertion “fruit is delicious,” with “apple is delicious” serving as the specific source. When someone encounters an unknown fruit, they can quickly understand its properties by associating it with the abstract knowledge of fruit, such as its possible taste or nutrition.
@@
Event Level: While a concept can capture the semantic meaning of a group of entities, it can also represent events at a higher level of conceptualization. Event-level conceptualization aims to broaden the scope from entities to include events as well He et al. (2024); Wang et al. (2024c). It seeks to associate different events under a shared concept that preserves the original semantic meaning to the maximum extent possible. For instance, activities like “Sam playing with his dog,” “Alex dancing in the club,” and “Bob doing yoga” can all be conceptualized as “relaxing events.” Abstract knowledge can then follow, stating that “If someone engages in relaxing events, they feel happy and relaxed.” When someone encounters an unknown or unfamiliar event, such as “Charlie likes painting the sunset,” they can infer that painting the sunset is a relaxing event and that Charlie feels happy and relaxed when doing so.
Document Level: Document-level conceptualization further extends the scope of the instance from entities and events to paragraphs or even entire documents. It aims to generate a summary that captures the main ideas and essential information while maintaining the overall semantic and context of the original text. Previous works on abstractive summarization Ladhak et al. (2022); Wang et al. (2019) have identical objectives, and earlier surveys by Rennard et al. (2023); Lin and Ng (2019); Liu et al. (2024) have effectively summarized these studies. Therefore, we only mention it here to clarify document-level conceptualization for readers and will not go into further detail in later sections to avoid overlap.
@@
System Level: Finally, system-level conceptualization aims to simplify the understanding of a complex system by abstracting its behavior and functionality into a higher-level representation. It is derived from the design of operating systems Doane et al. (1990) and is under-studied in the domain of NLP. The only representative example is a recent work by Subramonian et al. (2023), where the authors provide a systematic categorization of NLP tasks based on their objectives and characteristics while neglecting the detailed format of input/output and the datasets on which the tasks are evaluated. Due to the limited number of works available, we will not survey this type of conceptualization.
In later sections, we focus specifically on entity and event-level conceptualizations and propose a taxonomy to categorize works into three categories. To ensure that our search for papers is comprehensive and objective in relation to our target scope, we propose the following three objectives for selecting the most relevant papers. First, we aim for papers that adhere to the paradigm of linking different instances together and use concepts as representations of the formed clusters. We also seek papers that aim to establish hierarchies between different entities and events. Finally, we look for papers that directly seek abstractions of entities or events via concepts. Our proposed taxonomy is primarily categorized into resources, methods, and downstream applications of conceptualization, as this is the most straightforward structure for readers to grasp the topic.
@@ 3 Tasks and Datasets
We first survey available datasets and benchmarks, as well as their associated tasks, for these two types of conceptualizations. Statistical comparisons between different resources are shown in Table 1. For datasets that also serve as evaluation benchmarks, we mark their associated tasks with classification task (CLS) and generation task (GEN).
3.1 Concept Linking Task
Most conceptualizations can be formulated as a concept linking task, where the goal is to link an instance $i$ to a concept $c$ such that $i$ can be semantically represented by $c$. It is challenging due to the infinite number of possible instance-concept pairs. Previous approaches, such as those by Brauer et al. (2010); Yates et al. (2015), have attempted to further restrict the task to linking instances to a limited set of strict ontologies using heuristic or statistical methods. The task can also be formulated with a generative objective, which requires a model to generate $c$ directly given $i$ as input.
3.2 Entity-level Conceptualization Datasets
@@
3.3 Event-level Conceptualization Datasets
@@ 4 Conceptualization Acquisition Methods
Next, we survey methods for performing or collecting entity and event-level conceptualizations. We categorize them into three paradigms: extraction, retrieval, and generative-based methods, which are briefly demonstrated in Figure 3. We provide more discussions in Appendix A.
@@ 4.2.2 Neural-Based Retrieval
Neural-based retrieval methods overcome previous limitations by leveraging neural networks (or language models) to learn the semantic representations of the input instance and the concepts in the knowledge base or concept taxonomy. Then, the similarity between the input instance and the concepts can be calculated based on the learned representation embeddings. This approach can be benefitted by the advancement in language modeling, such as BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), and DeBERTa He et al. (2021, 2023). The most representative work in neural-based concept retrieval is AbstractATOMIC He et al. (2024). It uses GlossBERT Huang et al. (2019) to encode concepts (from WordNet and Probase) and instances (extracted from events in ATOMIC Sap et al. (2019)) into embeddings and leverage cosine similarity and human annotations to collect conceptualizations in a large scale manner. Other methods Wang et al. (2024c); Zhang et al. (2020, 2022); Lu et al. (2023); Wang et al. (2023b); Gao et al. (2022); Becker et al. (2021) similarly adopt different strategies in leveraging LMs as encoders, expanding the coverage of instances,training retrieval models. Despite their promising results, these methods are limited by their need for extensive labeled data, reliance on the completeness and accuracy of the knowledge base, and inability to retrieve new concepts that are out of training data.
@@ 4.3.1 Fine-Tuning-Based Generative Methods
Fine-tuning-based generative methods aim to take an entity or event as input and generate the concept directly via a fine-tuned generative language model. This approach allows the model to generate conceptualizations for new instances and offers maximum flexibility of the input. Several methods Peng et al. (2022); Yuan et al. (2023); He et al. (2024); Wang et al. (2024c, b, 2023b) have adopted this paradigm in training generative conceptualizers, based on models such as GPT2 Radford et al. (2019), BART Lewis et al. (2020), and T5 Raffel et al. (2020), for automated conceptualization acquisition. These methods typically train LMs on human-annotated or pre-existing conceptualization resources and yield outstanding results. However, fine-tuning-based generative methods are limited by their high computational cost, time-consuming and resource-intensive data collection, uncertain performance across diverse domains, and relatively low quality of novel concepts compared to human annotations Liu et al. (2025). While these are common limitations associated with fine-tuned generative models, zero-shot generative methods using powerful LLMs and advanced prompting techniques potentially address these issues.
@@ 5 Downstream Applications
We then survey downstream tasks that can benefit from applying conceptualizations to provide readers with a general picture of what can be achieved and how to benefit from integrating conceptualizations. An overview of performances by different methods that leverage conceptualization, evaluated on various benchmarks, are shown in Figure 4.
Figure 4: Empirical benefits of conceptualization in methods across various benchmarks compared to baselines.
@@ 5.1 Commonsense Reasoning
The task of generative commonsense inference modeling (COMET; Bosselut et al. (2019); Hwang et al. (2021)) aims to complete an inferential commonsense knowledge given a head event and a commonsense relation. State-of-the-art methods for COMET mainly fine-tune language models on large-scale commonsense knowledge bases, which suffer from data sparsity and lack of diversity in commonsense knowledge. Although transfer from LLMs helps West et al. (2022, 2023), distilled knowledge tends to be too easy for models to learn and converge to trivial inferences. To address these issues, Wang et al. (2023b) proposed to leverage conceptualization as knowledge augmentation tools to improve COMET. Conceptualizations are first derived from head events to obtain abstracted events. Then, the tail of the original commonsense knowledge is placed back to the abstracted event to form abstracted commonsense knowledge. These derived abstract knowledge are then integrated with the original knowledge in commonsense knowledge bases to enrich the diversity of commonsense knowledge. Experiments show consistent improvement in models’ performances. Wang et al. (2024a) further show that, by instantiating conceptualizations in abstract knowledge back to other novel instances, models can be further improved by training with newly instantiated knowledge. Liu et al. (2023a) also proposed a task that aims to generate diverse sentences describing concept relationships in various everyday scenarios. Conceptualizations and associated abstract knowledge can further boost models’ performances on this task.
@@ 6 Future Directions and Conclusions
6.1 Controllable Generation
@@ 6.2 Modeling Changes in Distribution
Conceptualization also plays a pivotal role in building reasoning systems that can capture situational changes in distribution to achieve System II reasoning Sloman (1996); Kahneman (2011). Among the several components that make up System II reasoning, a key element is the ability to reason with situational changes in distribution Bengio et al. (2021, 2019). These changes are triggered by environmental factors and actions by the agents themselves or others, especially when dealing with non-stationarities  Bengio (2017); Xu et al. (2025); Shi et al. (2025). This ability can be achieved by dynamically recombining existing concepts in the given environment or action and learning from the resultant situational changes Lake and Baroni (2018); Bahdanau et al. (2019); de Vries et al. (2019). For instance, consider the event “PersonX is driving a car on a sunny day.” A change in the weather from sunny to rainy could cause a different outcome, such as “PersonX becomes more cautious and drives slower.” This illustrates that a change in weather conditions can lead to a change in the driver’s behavior, representing an environmental change that triggers situational changes within the distribution of different weather conditions. In this process, the model is required to infer different changes that can possibly occur within a single event as the context, and reason about the potential outcome of each change. To model the distribution of different changes within an event, conceptualization can be used to represent the different states of the environment or action Wang and Song (2025). The model can then reason about the changes in distribution by manipulating the granularity of conceptualized changes. This type of distributional conceptualization not only provides an ontology for modeling the distribution of different changes within an event, but also assists the model in reasoning about the potential outcomes with appropriate abstract knowledge. Future works can leverage LLMs to curate benchmark datasets via sequential conceptualization generation and develop advanced systems for System II reasoning.